A ball traveling in the +x direction hits the wall at 30m/s and rebounds at 25m/s. If the ball is in contact with the wall for 4ms, determine the average acceleration of the ball during this time interval.

1
Expert's answer
2022-01-30T13:52:32-0500

The average acceleration of the ball during this time interval is simple change in speed divided by time it took:


"a=\dfrac{v_f-v_i}{t}=\dfrac{-25-30}{0.004}=13750\text{ m/s}^2."